Twelve more than a square of a number is 7 times the number. find the number.
Thank you

Answer by fractalier(2101)

Let the number be x.
12 + x^2 = 7x
x^2 - 7x + 12 = 0
(x - 3)(x - 4) = 0
x = 3 or x = 4
they both work.